OpenStudy (anonymous):

Is there a quick way to find the index of nilpotency of a matrix? Like A=[0 2 1; 0 0 3; 0 0 0]?

OpenStudy (anonymous):

without having to multiply the matrix by itself until we get all zeroes

OpenStudy (anonymous):

for a 3x3 this shouldnt be very difficult

OpenStudy (anonymous):

Yeah I know but I have exam tomorrow so just worried I might get a 4x4 or the power for the matrix might be high

OpenStudy (anonymous):

By index I assume the number of times it needs to be composed to destroy it?

OpenStudy (anonymous):

If thats the case then you just look at the lowest diagonal

OpenStudy (anonymous):

What do you mean by lowest diagonal?

OpenStudy (anonymous):

What do you mean by index?

OpenStudy (anonymous):

The smallest n such that A^n=0?

OpenStudy (anonymous):

like for this example when the matrix A is A^3 it's all zeroes so nilpotency index is 3

OpenStudy (anonymous):

Just look at the first diagonal with entries, and see how many diagonals including that one are above it. Then A^(that number + 1)

OpenStudy (anonymous):

oh so if there's only 1 entry that's not a diagonal?

OpenStudy (anonymous):

Suppose only the corner entry has a value. So that is only one diagonal, so its A^(1+1) or A^2=0

OpenStudy (anonymous):

Oh okay that make sense so then you'd know the index of a 3x3 is always 3

OpenStudy (anonymous):

No! it depends on the lowest diagonal with values.

OpenStudy (anonymous):

It could be either 3 or 2

OpenStudy (anonymous):

If only the corner is filled in a 3x3 matrix it would be 2

OpenStudy (anonymous):

okay I just want to make sure we're on the same page.. A=[ 0 0 0; 1 0 0; 8 1 0] it's 3 because of the 0 0 0 and 1 1 ?

OpenStudy (anonymous):

but if it's like all zeroes then nilpotency index can only be 2 max

OpenStudy (anonymous):

Index = 2 \[\left[\begin{matrix}0 & 0 & 1 \\ 0 & 0 & 0 \\ 0 & 0 & 0\end{matrix}\right]\] Index = 3 \[\left[\begin{matrix}0 & 1 & 0 \\ 0 & 0 & 1 \\ 0 & 0 & 0\end{matrix}\right]\]

OpenStudy (anonymous):

Okay that's very clear that's a lot!